OpenStudy (anonymous):

can someone please help only have 2 hours til deadline

OpenStudy (pinksapphire):

I can help you with the second one cause the first one won't open...

OpenStudy (anonymous):

ok thank you

OpenStudy (pinksapphire):

1.a)\[A=a^2\] a=the side \[A=18^2\] So what is A?

OpenStudy (anonymous):

i dont know i have that to but i don know where to go on from that

OpenStudy (pinksapphire):

What is\[18\times18\]

OpenStudy (pinksapphire):

\[18\times18=324\]

OpenStudy (anonymous):

so a =324?

OpenStudy (pinksapphire):

Yes.

OpenStudy (anonymous):

ok thank you can you help me with any of the other questions

OpenStudy (pinksapphire):

1.b)\[V = \frac{ bh }{ 3 }\]

OpenStudy (pinksapphire):

\[Volume=\frac{ 20\times18 }{ 3 }\]

OpenStudy (pinksapphire):

Volume=120

OpenStudy (anonymous):



OpenStudy (anonymous):

we have a right triangle so we can use our trig functions

OpenStudy (anonymous):

also we have 1 side and 2 angles so we can easily find the 3rd angle

OpenStudy (anonymous):

to find the missing angle we do \[\angle = 180 - 90 - 49 = 41 degrees\]

OpenStudy (anonymous):

Using that angle we can find the height of the plank using the function \[Cos \theta = Adjacent/ Hypotenuse\]

OpenStudy (anonymous):

\[Cos(41) = 3 / x\] From there we see our angle is 41 the adjacent side is the ground and the hypotenuse it the plank so now to find the side just solve for x

OpenStudy (anonymous):

To find the wall, using the same angle we would use\[Tan \theta = Opposite/Adjacent\] Seeing as our angle is 41 the opposite side of that is the wall and the adjacent side is the ground we now have \[Tan(41) = x/3\] now you just solve for x

OpenStudy (anonymous):

OK and then what

OpenStudy (anonymous):

Once you solve for both of those X values you will have your answers